OpenStudy (anonymous):

Find an equation of the line through A (-6, 5) having slope 7

OpenStudy (zepp):

In the slope-intercept form, plug (x,y) in to find the equation.

OpenStudy (henryblah):

General equation of a straight line. Either y=mx+c or (y-y1)=m(x-x1) Slope 7 therefore substitute 7 into the equation as m. Then put in x and y.

OpenStudy (zepp):

The slope-intercept being \[y=mx+b\]\[5=(7(-6)+b\]

OpenStudy (anonymous):

but in this question only one point given?

OpenStudy (zepp):

Yes, but you have the slope.

OpenStudy (zepp):

You only need a point and a slope to find a straight line function.

OpenStudy (anonymous):

so I use this formula for finding equation of line : y=mx+b? Or I can also use this one y-y1 = m (x-x1) ?

OpenStudy (henryblah):

Either is acceptable, but you generally use the 1st if you've just started doing this sort of stuff.

OpenStudy (zepp):

E.g.: I have a function g Its slope is 0 and passes by the origin  You plug everything in the equation to find the x-intercept \[f(g)=mx+b\]\[0=0x+b\]\[b=0\] \[f(g) = 0x + 0\]\[f(g)=x\] The formula \[\large m= \frac{y_{2} - y_{1}}{x_{2}-x_{1}},~~m=slope\] is usually used when you want to find the slope with 2 points given.

OpenStudy (anonymous):

Thanks zepp I only just need to plug the values or do simplification in it?

OpenStudy (zepp):

Yes, what I did above is to find the y-intercept, you already have the slope. Remember, a straight line function is defined by 2 things: the slope and the y-intercept

OpenStudy (anonymous):

So I don't need more simplication, that's my answer ?

OpenStudy (anonymous):

y=mx+b 5=7(-6)+b ?

OpenStudy (zepp):

Yes! ;D

OpenStudy (henryblah):

The reason why you use y-y1 etc is because that is the totally general formula. You will notice sonner or later that y=mx+c does not work for vertical lines etc. Also it is easier to manipulate in some circumstances

OpenStudy (zepp):

Find 'b'

OpenStudy (anonymous):

Thank you, and sorry my English is not good.

OpenStudy (amistre64):

my contribution is this: \[y-y_o=m(x-x_o)\] \[y=y_o=mx-mx_o\] \[y=mx-mx_o+y_o\] \[y = mx+b; \ b=y_o-mx_o\]

OpenStudy (amistre64):

got a spurious "=" in there

OpenStudy (zepp):

lol two equals

OpenStudy (amistre64):

:) the point is; you dont have to resolve for b, its already there

OpenStudy (anonymous):

it means b = 7(-6) +5

OpenStudy (anonymous):

and now if I solve it, I will find the interecpt ?

OpenStudy (amistre64):

\[y=mx-mx_o+y_o\] \[y=7x-7(-6)+5\]

OpenStudy (anonymous):

but y is given in this question?

OpenStudy (amistre64):

xo and yo are given; not x and y x and y are generic points for input and output of the equation itself xo and yo are points to establish the equation with and are constant

OpenStudy (zepp):

Thought it would be lot easier with the method I showed? D: \[y=mx+b\]\[5=7(-6)+b\]\[5=-42+b\]\[b = 5--42\]\[b=47\] Then plug everything back \[y=mx+b\]\[y =7(the slope)x + 48\]\[y=7x+47\]

OpenStudy (zepp):

got a spurious "8" in there

OpenStudy (amistre64):

those 8s can be tricky

OpenStudy (anonymous):

Thanks Zep, your procedure is simple compared to amistre but thank you both of you :)

OpenStudy (zepp):

 No offence, @amistre64, just a joke :P Np ;)